Charlie Chairs Inc., manufactures plastic moldings for car seats. Its costing system utilizes two cost categories, direct materials and conversion costs. Each product must pass through Department A and Department B. Direct materials are added at the beginning of production. Conversion costs are allocated evenly throughout production.

Data for Department A for February 2017 are:

WIP, beginning inventory, 40% completed               200 units

Units started during February 1000 units

WIP, ending inventory, 50% completed 240 units

Costs for Department A for Feburary 2015 are:

WIP, beginning inventory:                                        

Direct Materials                                                          $150,000

Conversion Costs                                                       208,000

Direct Materials costs added during February           606,000

Conversion costs added during February                  431,000

What is the unit cost per equivalent unit of beginning inventory in Department A?
Answer #1

Solution:

Equivalent unit of direct material in beginning inventory = 200*100% = 200

Equivalent unit of conversion in beginning inventory = 200*40% = 80

Cost per equivalent unit of material in beginning inventory = $150,000 / 200= $750 per unit

Cost per equivalent unit of conversion in beginning inventory = $208,000 / 80= $2,600 per unit

unit cost per equivalent unit of beginning inventory in Department A = $750 + $2,600 = $3,350 per unit